2. A retailer sells 10,000 units at $10. Each unit costs the retailer $7.00 to sell, with many parts imported from Japan. A marketing manager is rewarded on a revenue target & gets a bonus of 10% of revenue earned if total revenues meet/exceed $30k. The federal government has offered reduced tariffs for components of your product if you choose to import those parts from Mexico instead of Japan (where you currently get them). If you choose to import those parts from Mexico, you can reduce the price of your product to $7, while reducing costs to $3. A proposed price cut to $7 with reducing costs to $3 will boost sales up to 13,000 units. a.) Marketer’s opinion: Should you import from Mexico or Japan with corresponding price strategy? b.) Firms opinion: Should you import from Mexico or Japan with corresponding price strategy?
Answer #1

Current Plan (import from Japan)

Demand = 10000 units

Price = $10 per unit

Cost = $7 per unit

Revenue = 10*10000 = $100000

Bonus for Marketer = 10% of revenue = $10000

Profit to firm = revenue – cost – bonus = 100000 - $7*10000 – 10000 = $20000

Changed plan (Import from Mexico)

Demand = 13000 units

Price = $7

Cost = $3

Contribution = Price – Cost = 7 – 3 = $4

Hence Revenue = 7*13000 = $91000

Bonus for Marketer = 10% of revenue = $9100

Profit for Firm = contribution * unit – bonus = 13000*4 – 9100 = 52000 – 9100 = 41900

a. Marketer’s Opinion: As the bonus for the marketer has been reduced, hence marketer wont like the option of importing from Mexico.

b. Firm’s Opinion: After deducting the bonus for the marketer, the firm is having more profit when imported from Mexico than from Japan. Hence the Firm will choose Mexico over Japan.

IF YOU LIKE THE ANSWER, PLEASE GIVE AN UP-VOTE OR THUMB UP. THIS WILL ENCOURAGE ME TO ANSWER MORE!!